Let f(x) = 3x and g(x) = f(x + 1) + 1 . If g(x) = 13 what is the value of x?

(A) 1
(B) 2
(C) 3
(D) 4
(E) 5

Answer by ikleyn(52794)

.

Equation for x is

    3(x+1) + 1 = 13.


From this equation, find x

    3(x+1) = 13-1 = 12

    x + 1 = 12/3 = 4

    x = 4-1 = 3.


ANSWER.  x = 3.

Solved.